

********************************************************************
UARS WEEKLY STATUS REPORT - 11 OCT 2002
********************************************************************
***********
SOLSTICE
***********
Nothing received.
***********
SUSIM
***********
SUSIM successfully performed its daily mid-resolution and MgII
solar scans during periods of predicted UARS telemetry.
Today 10/11/02 marks the 11th year of SUSIM's observations, SUSIM has now
collected data for an entire solar cycle.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
PEM continues to work with the GUVI team on the diagnosis of data from
the first conjunction period. Analysis of the two conjunction passes showed
that the aurora as detected by GUVI was at higher latitude than UARS.
GUVI images from other times from the first conjunction period show spatial
regions that both GUVI and PEM sampled when aurora was present. Separation
is several minutes to 10's of minutes when both instruments saw aurora. A
presentation will be made at the November American Meteorological Society
meeting on PEM's view of the first conjunction period. PEM has seen no aurora
during the second conjunction period.
PEM data for a third GUVI-PEM conjunction has been planned. This time
period is October 27. Orbit analysis shows that conjunctions will occur
between TIMED and UARS at 11:03 UT (IL = 65 deg, MLT = 20 hr, GeoLat = -51 deg,
GeoLon = 132 deg) and 11:50 UT (IL = 61 deg, MLT = 8 hr, GeoLat = 52 deg,
GeoLon = 295 deg).
PEM continues to receive NMC data files from NOAA. No problems have
been encountered. NOAA has changed the binary structure of the NMC data and
PEM is now receiving two data streams, one in the original VAX format and the
other from their new binary format.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact for UARS
instrument operations to support re-engineering of the UARS
Mission Operations Concept to develop capabilities for the
Science Traceability Mission.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 28 September 2002 (Orbit 60414, GMT Day 271)
through 04 October 2002 (Orbit 60534, GMT Day 277).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
PEM 272/1303 60437 POWER ON AXIS 1
HRDI 276/1835 60500 TO DAY ONLY MODE
The beta angle ranged between the angles of 43.6 degrees to 27.1
degrees for this report period. The beta angle is now decreasing
towards a minimum angle of 0 degrees on 10 October 2002 (DOY 283).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 0.96 to 0.97 degrees C. Additional
battery performance data for each battery on 28 September and 04
October 2002 are provided in Appendix D.
The clock error ranged from -12.6 to 10.4 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 04
PEM = 01
HRDI = 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 00
Two (2) Anomaly Report remained OPEN at the end of this report period
(see Appendix E).
GROUND SYSTEM OPERATION
Zero (0) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The new hard disk on the prototype was installed this week. The new
build of the Eclipse software has also been installed and can now be
tested. Once the testing is completed the software can be released
for operations. This will most likely not take place until the ERBS
integration is complete.
The FOT is also working hard in preparation for the college students
that will be in the control center to learn about spacecraft
operations. The FOT will try to formulate training materials and
structure a learning plan to teach the students. This is all part of
the new Educational Outreach initiative/contract, which started on
October 1, 2002 for both UARS and TOMS.
MEETINGS
- The FOT attended the UARS Monthly Status review on 03 October
2002.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 10 October 2002 (DOY 283).
This maneuver will be a Forward to Reverse Flight maneuver (Northern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 04 October 2002
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
60437 272/1303 02-340 PEM POWER ON AXIS 1
60475 275/0237 02-341 OBC/CLK CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
60494 276/0820 02-342 OBC/CLK CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
60500 276/1835 02-343 HRDI TO DAY-ONLY MODE
60501 276/2016 02-344 OBC/CLK CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
60519 277/2351 02-345 OBC/CLK CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 04 October 2002
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 04 October 2002
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
60475 275/0241 CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
60494 276/0822 CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
60501 276/2016 CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
60519 277/2353 CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 04 October 2002
27 September 2002 - GMT Day 270 Beta = 49.7 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -145.6/-11.2 -3.68/-4.63 +00.0/+00.0 08.3 0.000 0.0
2 +11.2/-50.4 +5.14/+3.23 +24.4/-15.2 25.0 1.030 19.8
3 +22.4/-33.6 +3.87/+2.60 +24.0/-15.2 25.0 1.040 19.3
04 October 2002 - GMT Day 277 Beta = 27.1 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -140.0/-168.0 -3.05/-4.31 +00.0/+00.0 08.3 0.000 0.0
2 +11.2/-89.6 +5.15/+4.19 +28.8/-15.2 24.8 1.030 21.0
3 +22.4/-67.2 +4.19/+3.23 +28.0/-15.2 24.8 1.020 20.5
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
-None this report period.
APPENDIX E
UARS ANOMALY REPORTS
Week ending 04 October 2002
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
02-050 SOLSTICE 60336 265/2037 N/A SOLS UNEXPECTED STATE
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
02-051 FOT 60366 267/1951 30832 GENERIC LATE ACQUISITION
02-052 SOLSTICE 60366 267/1951 N/A LOADING PROBLEM